# Unraveling time-varying causal effects of multiple exposures: integrating Functional Data Analysis with Multivariable Mendelian Randomization

### v2

@@ 1 Introduction
Establishing causal relationships between exposures and health outcomes remains a central challenge in epidemiological research. Mendelian Randomization (MR) is a widely used method for causal inference that leverages genetic variants as instrumental variables (IVs) to estimate the causal effect of an exposure on an outcome [sanderson2022mendelian, Lawlor2008]. Traditional MR methods generally assume that a single exposure exerts a constant causal effect. Since genetic variants are fixed at conception, estimates from traditional MR are typically interpreted as reflecting the average effect over a lifetime of genetically predicted exposure across the entire life course, reflecting cumulative rather than age-specific or period-specific influences [burgess2021interpretation]. However, many exposures exhibit time-varying effects, meaning that their causal impact on outcomes can change across different stages of life [Kaneko2023]. For example, the effect of body mass index on cardiovascular disease risk may differ between early adulthood and older age. This variation can occur due to evolving physiological processes or age-specific biological mechanisms [Palmer2022, Karlsson2020]. To capture these dynamics, analytical methods must move beyond single time-point estimates and instead model how effects evolve over time. This approach allows for the identification of age-specific or life course-specific influences that may be overlooked when relying on a single measurement of exposures [wagner2021time, shi2022mr_timevarying]. Moreover, emerging evidence also indicates that genetic associations with certain exposures can vary at different time points in the life course (within a population), meaning that MR estimates may reflect the effect of exposure at particular life periods rather than uniformly over time [sanderson2022timevarying, jiang2021age]. Accounting for these time-varying genetic effects allows researchers to identify critical periods when an exposure has the greatest impact on health outcomes [Power2024Mendelian]. Incorporating these dynamics into MR analyses, therefore, enhances causal inference for time-dependent exposures and can inform the design and timing of preventive interventions. In response to this challenge, recent work has focused on estimating time-varying causal effects using MR. Most existing methods discretize time, treating exposures measured at fixed points as separate variables, ignoring the continuous nature of exposure trajectories. Tian et al. [Tian2024] proposed a continuous-time MR framework, called Univariable Functional Mendelian Randomization (U-FMR), which models the exposure effect as a smooth function of time using functional data analysis [ramsay2005fda]. Their method uses Functional Principal Component Analysis (FPCA) to reduce the dimensionality of longitudinal exposure data and integrates these low-dimensional representations into a MR model. This framework provides a flexible and theoretically robust approach for inferring how causal effects evolve over the life course, given longitudinal data.
@@
Despite these advances, existing MR methods that investigate the life course effects of exposure have focused exclusively on single exposures [Tian2024, cao2016timevarying, richardson2020earlylater, shi2021timevarying, sanderson2022timevarying, richardson2023timevarying]. However, in real-world scenarios, multiple exposures often act simultaneously or influence each other over the life course, particularly when they are biologically related. Multivariable MR (MV-MR) extends the univariable MR (U-MR) framework to jointly estimate the causal effects of multiple exposures using a set of genetic variants as instruments [sanderson2019examination]. Unlike U-MR, which captures only the total effect of a single exposure, MV-MR can disentangle the direct effects of many exposures, provided that the genetic instruments are sufficiently associated with each exposure throughout the life course. However, existing MV-MR methods assume that causal effects are constant over time.
@@
To address the limitation of time-invariant effects, we propose Multivariable Functional Mendelian Randomization (MV-FMR), which extends U-FMR [Tian2024] to simultaneously model multiple time-varying exposures. Our key methodological contributions include: (1) a data-driven cross-validation procedure for selecting the number of basis components for each exposure, (2) joint modeling of time-varying exposures that share common genetic pathways, (3) handling of mediation pathways between time-varying exposures, and (4) implementation for both continuous and binary outcomes. To evaluate the performance of MV-FMR, we conducted extensive simulation studies under three main scenarios: horizontal pleiotropy through overlapping genetic instruments, mediation pathways where one exposure influences another, and null effects to assess specificity. We further performed sensitivity analyses examining different functional forms for the exposure-outcome effects (linear and nonlinear effects) and outcome types (continuous and binary). Results demonstrated that MV-FMR accurately recovered time-varying causal effects, consistently outperforming univariable functional MR approaches. We apply MV-FMR to UK Biobank data to estimate the time-varying causal effects of systolic blood pressure and body mass index on coronary artery disease risk across adulthood, demonstrating how the method can identify critical age periods and disentangle direct effects from mediation pathways that would be missed by traditional approaches.
@@ 2.1 Background on Multivariable Mendelian Randomization (MV-MR)
Multivariable MR (MV-MR) generalizes the classical MR framework to jointly estimate the direct causal effects of multiple exposures on an outcome using genetic variants as instrumental variables [sanderson2021multivariable]. While univariable MR captures the total effect of a single exposure, MV-MR disentangles the direct effect of each exposure while accounting for the influence of the others. Estimating multiple exposures simultaneously is often desirable, for example, when exposures are biologically or clinically related, or when one exposure may mediate the effect of another on the outcome. By modeling multiple exposures simultaneously, MV-MR can disentangle overlapping causal mechanisms and improve robustness to pleiotropy, provided that pleiotropic effects operate only through the modeled exposures. The validity of MV-MR requires an extension of the standard IV assumptions to the multivariable setting:
@@ 2.2 Overview of the Univariate Functional Mendelian Randomization (U-FMR) with continuous outcome
We begin by briefly showing the U-FMR model proposed by [Tian2024] as a foundation for our multivariable extension presented in 2.3. The setting involves a continuous outcome $Y$, a time-varying exposure $X(t)$, and a set $P$ of genetic instruments $\bm{G}$. Inference on time-varying causal effects relies on life-course variation in the associations between the genetic instruments and the exposure. When such associations evolve over time, they provide information that enables the identification of how the causal effect of the exposure on the outcome changes across the life course. Under this setting, the instruments act as valid proxies for the exposure across the observation period, enabling the estimation of the time-varying causal effect to be estimated. The causal effect is therefore modeled as a smooth function $\beta(t)$, capturing how the impact of $X(t)$ on $Y$ evolves over the observed time interval. This leads to the structural model:
@@
The coefficients $\beta^{*}_{k}$ are treated as pseudo-exposures, and estimated using the continuously updating Generalized Method of Moments (GMM) [hansen1996finite], where genetic variants $\bm{G}$ are used as instruments for the FPCA scores $\xi_{k}$. GMM is particularly well-suited for this setting as it performs robustly under weak instrument conditions commonly encountered with genetic instrumental variables. All variables are mean-centered prior to estimation to ensure numerical stability. The sample moment function is defined as:
@@ 2.3 Multivariable Functional Mendelian Randomization (MV-FMR) with continuous outcome
$=\underbrace{\beta_{0}+\int_{0}^{T}\left[\beta_{1}(t)\mu_{1}(t)+\beta_{2}(t)\mu_{2}(t)\right]dt}_{\beta_{0}^{*}}+\sum_{k_{1}=1}^{K_{1}}\underbrace{\left[\int_{0}^{T}\beta_{1}(t)\phi_{1,k_{1}}(t)\,dt\right]}_{\beta_{k_{1}}^{*}}\xi_{1,k_{1}}+\sum_{k_{2}=1}^{K_{2}}\underbrace{\left[\int_{0}^{T}\beta_{2}(t)\phi_{2,k_{2}}(t)\,dt\right]}_{\beta_{k_{2}}^{*}}\xi_{i2,k_{2}}+\tilde{\epsilon}$,
@@ 2.3.1 Estimation of component effects
Each exposure trajectory is represented using FPCA, with scores estimated via the PACE method [Yao2005] to accommodate sparse and irregular measurements. Full mathematical details of the FPCA representation and PACE estimation are provided in Supplementary Method 1. Let $\mathbf{G}=[\mathbf{G}_{1},\mathbf{G}_{2},\mathbf{G}_{12}]\in\mathbb{R}^{N\times P}$ denote the stacked matrix of genetic instruments, where: $\mathbf{G}_{j}\in\mathbb{R}^{N\times P_{j}}$ are instruments specific to $X_{j}(t)$, and $\mathbf{G}_{12}\in\mathbb{R}^{N\times P_{12}}$ are pleiotropic instruments influencing both $X_{1}(t)$ and $X_{2}(t)$. The estimation of $\boldsymbol{\beta}^{*}_{1}$ and $\boldsymbol{\beta}^{*}_{2}$ follows the GMM framework described in Section 2.2, extending it to jointly model both exposures. The stacked FPCA scores $\boldsymbol{\xi}=[\boldsymbol{\xi}_{1},\boldsymbol{\xi}_{2}]\in\mathbb{R}^{N\times(K_{1}+K_{2})}$ and coefficient vector $\boldsymbol{\beta}^{*}=[\boldsymbol{\beta}^{*}_{1},\boldsymbol{\beta}^{*}_{2}]$ are used in the GMM objective function with $\mathbf{G}$ as the full instrument matrix. This joint estimation allows the model to account for correlation between exposures through overlapping instruments while maintaining the asymptotic properties established for the univariable case.
@@ 2.3.3 Reconstructing the coefficient functions
where $\boldsymbol{\phi}_{j}(t)=[\phi_{j,1}(t),\ldots,\phi_{j,K_{j}}(t)]^{T}$ denotes the vector of eigenfunctions and $\boldsymbol{\hat{\beta}}^{*}_{j}\in\mathbb{R}^{K_{j}}$ is the corresponding vector of estimated coefficients.
@@ 2.4 Multivariable Functional Mendelian Randomization (MV-FMR) with binary outcome
For binary outcomes, we apply a two-stage residual inclusion (2SRI) control function approach to account for endogeneity in the FPCA scores. This method is preferred for nonlinear models over standard two-stage predictor substitution (2SPS) approaches, which replace endogenous variables with their predicted values. In the context of nonlinear models such as logistic regression, the 2SPS estimator produces inconsistent estimates because the nonlinear link function invalidates the substitution of predicted values, thereby distorting the conditional mean structure [terza2008two, wooldridge2015control]. In contrast, the 2SRI approach controls for endogeneity by including first-stage residuals as additional regressors. Following the notation introduced in 2.3, in the first stage, FPCA scores are regressed on instruments:
@@ 3.1 Simulation scenarios
  • Scenario 2: Null effect control (Figure 2, center left): Same genetic architecture as in Scenario 1, but the second exposure does not have a causal effect on the outcome $(\beta_{2}(t)=0\forall t)$. This scenario evaluates the type I error, ensuring that MV-FMR does not detect spurious causal effects.
  • Scenario 3: Mediation model (Figure 2, right left): $X_{1}(t)$ causally influences $X_{2}(t)$, and both affect the outcome $Y$. This scenario evaluates the ability of MV-FMR to recover direct effects while accounting for mediation.
@@
These three scenarios reflect common epidemiological challenges that arise when investigating multiple correlated risk factors [sanderson2022timevarying]. Horizontal pleiotropy is frequently observed when genetic variants influence multiple physiological pathways, potentially violating the exclusion restriction in univariable MR analysis. Mediation pathways are pervasive in chronic disease epidemiology, where one risk factor operates both directly on the outcome and indirectly through downstream exposures. In such cases, U-FMR estimates the total causal effect, whereas MV-FMR recovers the direct effect by conditioning on the mediator. Finally, it is important to assess specificity in settings where some exposures have true null effects. In practice, analyses of joint time-varying exposures may include variables that are correlated with causal risk factors but do not themselves affect disease risk, and we aim to verify that the model does not introduce spurious non-null estimates. Correctly distinguishing these relationships is essential for identifying modifiable targets for intervention and avoiding spurious causal inferences.
@@ 3.2 Data generating process
Functional principal components are calculated using the fdapace package [fdapace]. We compare the performance of MV-FMR against U-FMR applied separately to each exposure. Performance is assessed using the Integrated Squared Error (ISE), which quantifies the average squared deviation between estimated and true effect functions, and the pointwise 95% coverage rate, defined as the proportion of time points where the 95% confidence intervals contain the true effect. Each simulation scenario is replicated $R=500$ times with a sample size $N=5000$. The number of instruments is set to $P_{1}=P_{2}=100-P12$, where $P_{12}=p_{12}\cdot 100$ with $p_{12}=0.15$ in the baseline simulations. Additional simulations vary the degree of instrument overlap across exposures, considering values of $p_{12}$ equal to 0.15, 0.3, 0.5 and 0.8, to explore how different levels of overlap affect MV-FMR performance. Instrument strength is measured using the conditional F-statistic for each principal component [Sanderson2021-Fstat].
@@ 3.3 Simulation results
We then evaluated the robustness of MV-FMR to different functional forms of the exposure-outcome relationships for Scenarios 1 and 3. Specifically, we considered three distinct combinations: (i) both exposures with linear effects, (ii) one exposure with a linear effect and the other with a quadratic effect, and (iii) both exposures with quadratic effects. Figure 3 presents the estimated coefficient functions $\beta_{1}(t)$ and $\beta_{2}(t)$ averaged across simulations for a continuous outcome, with corresponding results for a binary outcome provided in Supplementary Figure 1. Across all time points, MV-FMR consistently achieved lower mean squared error (MSE) than the univariable approach and more accurately recovered the true functional forms. Notably, MV-FMR successfully captured nonlinear time-varying effects, whereas the univariable failed to reproduce the underlying functional relationships. Additional simulations examined alternative exposure–outcome functional forms, including stepwise and decreasing lifetime effects (Supplementary Table 3). Together, these results confirm the robustness of MV-FMR across a wider range of true effect shapes.
@@
We applied the proposed MV-FMR framework to investigate the time-varying causal effects of systolic blood pressure (SBP) and body mass index (BMI) on coronary artery disease (CAD) risk. This application provides a good case study for our functional extension of multivariable MR for several reasons. First, SBP and BMI are well-established cardiovascular risk factors with substantial phenotypic correlation [hall2015obesity, collaboration2016body], creating a scenario in which multivariable approaches offer advantages over univariable methods. Second, biological evidence indicates that BMI influences cardiovascular outcomes both directly and indirectly through blood pressure [seravalle2014obesity], establishing a mediation pathway that traditional univariable MR cannot disentangle. Indeed, cross-sectional multivariable MR studies have shown that approximately 27% of the BMI-CAD association operates through SBP [Gill2021], providing a well-characterized benchmark for validation. Third, both exposures exhibit time-varying effects over the lifecourse, with evidence suggesting that critical exposure windows for cardiovascular risk may differ between midlife and later life [Gnatiuc2017, duncan2021differential]. By jointly modeling SBP and BMI trajectories while accounting for their correlation and potential mediation, MV-FMR enables estimation of the direct time-varying causal effect of each exposure on CAD risk, allowing us to recover established causal patterns and uncover critical windows of effect that cannot be identified using standard cross-sectional or univariable approaches.
@@
To address this question, we used data from UK Biobank, a large-scale prospective cohort study that recruited over 500,000 participants aged 40-69 years between 2006 and 2010 across the United Kingdom [Sudlow2015]. For this analysis, we restricted to participants who met the following criteria: (i) European ancestry, (ii) availability of genetic data, and (iii) at least two longitudinal measurements of both SBP and BMI between ages 50 and 70 years. After applying these criteria, 59,907 participants were included in the final sample. SBP and BMI trajectories were reconstructed from all available measurements in primary care records and UK Biobank baseline assessment visits within the 50-70 year window. Incident CAD was defined using first-occurrence ICD-10 codes I21-I25, covering acute myocardial infarction and chronic ischemic heart disease.
We investigated the effects of SBP and BMI on CAD using 274 and 199 genetic variants, respectively. Genetic instruments were obtained from large-scale genome-wide association study (GWAS) meta-analyses: for SBP were taken from Evangelou et al. [Evangelou2018], and for BMI from Winkler et al. [Winkler2015]. Five variants were associated with both exposures, representing overlapping instruments that may introduce pleiotropy between SBP and BMI. We then applied MV-FMR to jointly estimate the time-varying causal effects of SBP and BMI on CAD risk over ages 50-70 years. Individual exposure trajectories were modeled using FPCA with PACE, with the number of components chosen via our data-driven k-fold cross-validation procedure. Two principal components were selected for both exposures. For SBP, the first two components explained 77.0% and 16.5% of the variability, respectively, while for BMI, they explained 90.5% and 8.1%. The large proportion of explained variance indicates that trajectories of both exposures were well captured by the two-component FPCA representation.
@@
Instrument strength was assessed using F-statistics for each principal component. For SBP, the first and second components yielded conditional F-statistics of 9.77 and 3.44, respectively, while for BMI, the corresponding values were 9.03 and 2.57. Although components fell below the conventional threshold for strong instruments ($F<10$), our simulation results indicate that functional MR approaches remain unbiased in such settings. To assess the benefits of joint modeling, we compared MV-FMR with U-FMR applied separately to each exposure. This comparison allowed us to evaluate whether simultaneous modeling of SBP and BMI revealed different causal effect patterns compared to the univariable analyses. Figure 4 presents the MV-FMR estimates. The analysis revealed distinct time-varying causal effects of SBP and BMI on CAD risk. Both exposures showed positive associations with CAD risk between ages 50 and 60, but their effects attenuated thereafter, becoming non-significant (95% bootstrap CI) beyond age 60. These findings suggest that the critical exposure window for SBP and BMI in relation to CAD risk is concentrated in the sixth decade of life.
@@
The comparison between MV-FMR and U-FMR provides evidence consistent with established mediation patterns. For SBP, the multivariable and univariable analyses yielded broadly consistent results, both identifying similar effect magnitudes and time windows of significance. For BMI, MV-FMR estimates showed attenuation in the point estimates compared to the univariable approach, suggesting that part of BMI’s apparent effect on CAD operates indirectly through blood pressure pathways. Inspection of the confidence intervals reveals partial overlap between MV-FMR and U-FMR estimates for BMI. Notably, this overlapping is also observed in the traditional MV-MR analysis [Gill2021], where the univariable BMI effect (OR 1.49, 95% CI 1.39 to 1.60) and the multivariable effect (OR 1.34, 95% CI 1.24 to 1.45) showed overlapping confidence intervals.

### v3

@@ 1 Introduction
Establishing causal relationships between exposures and health outcomes remains a central challenge in epidemiological research. Mendelian Randomization (MR) is a widely used method for causal inference that leverages genetic variants as instrumental variables (IVs) to estimate the causal effect of an exposure on an outcome [19, 12]. Traditional MR methods generally assume that a single exposure exerts a constant causal effect. Since genetic variants are fixed at conception, estimates from traditional MR are typically interpreted as reflecting the average effect over a lifetime of genetically predicted exposure across the entire life course, reflecting cumulative rather than age-specific or period-specific influences [1]. However, many exposures exhibit time-varying effects, meaning that their causal impact on outcomes can change across different stages of life [13]. For example, the effect of body mass index on cardiovascular disease risk may differ between early adulthood and older age. This variation can occur due to evolving physiological processes or age-specific biological mechanisms [14, 11]. To capture these dynamics, analytical methods must move beyond single time-point estimates and instead model how effects evolve over time. This approach allows for the identification of age-specific or life course-specific influences that may be overlooked when relying on a single measurement of exposures [30, 26]. Moreover, emerging evidence also indicates that genetic associations with certain exposures can vary at different time points in the life course (within a population), meaning that MR estimates may reflect the effect of exposure at particular life periods rather than uniformly over time [20, 10]. Accounting for these time-varying genetic effects allows researchers to identify critical periods when an exposure has the greatest impact on health outcomes [15]. Incorporating these dynamics into MR analyses, therefore, enhances causal inference for time-dependent exposures and can inform the design and timing of preventive interventions. In response to this challenge, recent work has focused on estimating time-varying causal effects using MR. Most existing methods discretize time, treating exposures measured at fixed points as separate variables, ignoring the continuous nature of exposure trajectories. Tian et al. [29] proposed a continuous-time MR framework, called Univariable Functional Mendelian Randomization (U-FMR), which models the exposure effect as a smooth function of time using functional data analysis [16]. Their method uses Functional Principal Component Analysis (FPCA) to reduce the dimensionality of longitudinal exposure data and integrates these low-dimensional representations into a MR model. This framework provides a flexible and theoretically robust approach for inferring how causal effects evolve over the life course, given longitudinal data.
@@
Despite these advances, existing MR methods that investigate the life course effects of exposure have focused exclusively on single exposures [29, 2, 18, 25, 20, 17]. However, in real-world scenarios, multiple exposures often act simultaneously or influence each other over the life course, particularly when they are biologically related. Multivariable MR (MV-MR) extends the univariable MR (U-MR) framework to jointly estimate the causal effects of multiple exposures using a set of genetic variants as instruments [21]. Unlike U-MR, which captures only the total effect of a single exposure, MV-MR can disentangle the direct effects of many exposures, provided that the genetic instruments are sufficiently associated with each exposure throughout the life course. However, existing MV-MR methods assume that causal effects are constant over time.
@@
To address the limitation of time-invariant effects, we propose Multivariable Functional Mendelian Randomization (MV-FMR), which extends U-FMR [29] to simultaneously model multiple time-varying exposures. Our key methodological contributions include: (1) a data-driven cross-validation procedure for selecting the number of basis components for each exposure, (2) joint modeling of time-varying exposures that share common genetic pathways, (3) handling of mediation pathways between time-varying exposures, and (4) implementation for both continuous and binary outcomes. To evaluate the performance of MV-FMR, we conducted extensive simulation studies under three main scenarios: horizontal pleiotropy through overlapping genetic instruments, mediation pathways where one exposure influences another, and null effects to assess specificity. We further performed sensitivity analyses examining different functional forms for the exposure-outcome effects (linear and nonlinear effects) and outcome types (continuous and binary). Results demonstrated that MV-FMR accurately recovered time-varying causal effects, consistently outperforming univariable functional MR approaches. We apply MV-FMR to UK Biobank data to estimate the time-varying causal effects of systolic blood pressure and body mass index on coronary artery disease risk across adulthood, demonstrating how the method can identify critical age periods and disentangle direct effects from mediation pathways that would be missed by traditional approaches.
@@ 2.1 Background on Multivariable Mendelian Randomization (MV-MR)
Multivariable MR (MV-MR) generalizes the classical MR framework to jointly estimate the direct causal effects of multiple exposures on an outcome using genetic variants as instrumental variables [23]. While univariable MR captures the total effect of a single exposure, MV-MR disentangles the direct effect of each exposure while accounting for the influence of the others. Estimating multiple exposures simultaneously is often desirable, for example, when exposures are biologically or clinically related, or when one exposure may mediate the effect of another on the outcome. By modeling multiple exposures simultaneously, MV-MR can disentangle overlapping causal mechanisms and improve robustness to pleiotropy, provided that pleiotropic effects operate only through the modeled exposures. The validity of MV-MR requires an extension of the standard IV assumptions to the multivariable setting:
@@ 2.2 Overview of the Univariate Functional Mendelian Randomization (U-FMR) with continuous outcome
We begin by briefly showing the U-FMR model proposed by [29] as a foundation for our multivariable extension presented in 2.3. The setting involves a continuous outcome $Y$, a time-varying exposure $X(t)$, and a set $P$ of genetic instruments $\bm{G}$. Inference on time-varying causal effects relies on life-course variation in the associations between the genetic instruments and the exposure. When such associations evolve over time, they provide information that enables the identification of how the causal effect of the exposure on the outcome changes across the life course. Under this setting, the instruments act as valid proxies for the exposure across the observation period, enabling the estimation of the time-varying causal effect to be estimated. The causal effect is therefore modeled as a smooth function $\beta(t)$, capturing how the impact of $X(t)$ on $Y$ evolves over the observed time interval. This leads to the structural model:
@@
The coefficients $\beta^{*}_{k}$ are treated as pseudo-exposures, and estimated using the continuously updating Generalized Method of Moments (GMM) [9], where genetic variants $\bm{G}$ are used as instruments for the FPCA scores $\xi_{k}$. GMM is particularly well-suited for this setting as it performs robustly under weak instrument conditions commonly encountered with genetic instrumental variables. All variables are mean-centered prior to estimation to ensure numerical stability. The sample moment function is defined as:
@@ 2.3 Multivariable Functional Mendelian Randomization (MV-FMR) with continuous outcome
$=\underbrace{\beta_{0}+\int_{0}^{T}\left[\beta_{1}(t)\mu_{1}(t)+\beta_{2}(t)\mu_{2}(t)\right]dt}_{\beta_{0}^{*}}+\sum_{k_{1}=1}^{K_{1}}\underbrace{\left[\int_{0}^{T}\beta_{1}(t)\phi_{1,k_{1}}(t)\,dt\right]}_{\beta_{k_{1}}^{*}}\xi_{1,k_{1}}+\sum_{k_{2}=1}^{K_{2}}\underbrace{\left[\int_{0}^{T}\beta_{2}(t)\phi_{2,k_{2}}(t)\,dt\right]}_{\beta_{k_{2}}^{*}}\xi_{2,k_{2}}+\tilde{\epsilon}$,
@@ 2.3.1 Estimation of component effects
Each exposure trajectory is represented using FPCA, with scores estimated via the PACE method [33] to accommodate sparse and irregular measurements. Full mathematical details of the FPCA representation and PACE estimation are provided in Supplementary Method 1. Let $\mathbf{G}=[\mathbf{G}_{1},\mathbf{G}_{2},\mathbf{G}_{12}]\in\mathbb{R}^{N\times P}$ denote the stacked matrix of genetic instruments, where: $\mathbf{G}_{j}\in\mathbb{R}^{N\times P_{j}}$ are instruments specific to $X_{j}(t)$, and $\mathbf{G}_{12}\in\mathbb{R}^{N\times P_{12}}$ are pleiotropic instruments influencing both $X_{1}(t)$ and $X_{2}(t)$. The estimation of $\boldsymbol{\beta}^{*}_{1}$ and $\boldsymbol{\beta}^{*}_{2}$ follows the GMM framework described in Section 2.2, extending it to jointly model both exposures. The stacked FPCA scores $\boldsymbol{\xi}=[\boldsymbol{\xi}_{1},\boldsymbol{\xi}_{2}]\in\mathbb{R}^{N\times(K_{1}+K_{2})}$ and coefficient vector $\boldsymbol{\beta}^{*}=[\boldsymbol{\beta}^{*}_{1},\boldsymbol{\beta}^{*}_{2}]$ are used in the GMM objective function with $\mathbf{G}$ as the full instrument matrix. This joint estimation allows the model to account for correlation between exposures through overlapping instruments while maintaining the asymptotic properties established for the univariable case.
@@ 2.3.3 Reconstructing the coefficient functions
where $\boldsymbol{\phi}_{j}(t)=[\phi_{j,1}(t),\ldots,\phi_{j,K_{j}}(t)]^{T}$ denotes the vector of eigenfunctions and $\boldsymbol{\hat{\beta}}^{*}_{j}\in\mathbb{R}^{K_{j}}$ is the corresponding vector of estimated coefficients. We employ a non-parametric bootstrap procedure for constructing confidence intervals. By resampling individuals with replacement, we construct pointwise 95% confidence intervals for $\hat{\beta}_{j}(t)$ using the $2.5^{\text{th}}$ and $97.5^{\text{th}}$ percentiles of the bootstrap distribution at each time point $t$. This procedure accounts for the uncertainty propagated through all stages of the estimation process. Full details are provided in Supplementary Method 2.
@@ 2.4 Multivariable Functional Mendelian Randomization (MV-FMR) with binary outcome
For binary outcomes, we apply a two-stage residual inclusion (2SRI) control function approach to account for endogeneity in the FPCA scores. This method is preferred for nonlinear models over standard two-stage predictor substitution (2SPS) approaches, which replace endogenous variables with their predicted values. In the context of nonlinear models such as logistic regression, the 2SPS estimator produces inconsistent estimates because the nonlinear link function invalidates the substitution of predicted values, thereby distorting the conditional mean structure [28, 32]. In contrast, the 2SRI approach controls for endogeneity by including first-stage residuals as additional regressors. Following the notation introduced in 2.3, in the first stage, FPCA scores are regressed on instruments:
@@ 3.1 Simulation scenarios
  • Scenario 2: Null effect control (Figure 2, center panel): Same genetic architecture as in Scenario 1, but the second exposure does not have a causal effect on the outcome $(\beta_{2}(t)=0\forall t)$. This scenario evaluates the type I error, ensuring that MV-FMR does not detect spurious causal effects.
  • Scenario 3: Mediation model (Figure 2, right panel): $X_{1}(t)$ causally influences $X_{2}(t)$, and both affect the outcome $Y$. This scenario evaluates the ability of MV-FMR to recover direct effects while accounting for mediation.
@@
These three scenarios reflect common epidemiological challenges that arise when investigating multiple correlated risk factors [20]. Horizontal pleiotropy is frequently observed when genetic variants influence multiple physiological pathways, potentially violating the exclusion restriction in univariable MR analysis. Mediation pathways are pervasive in chronic disease epidemiology, where one risk factor operates both directly on the outcome and indirectly through downstream exposures. In such cases, U-FMR estimates the total causal effect, whereas MV-FMR recovers the direct effect by conditioning on the mediator. Finally, it is important to assess specificity in settings where some exposures have true null effects. In practice, analyses of joint time-varying exposures may include variables that are correlated with causal risk factors but do not themselves affect disease risk, and we aim to verify that the model does not introduce spurious non-null estimates. Correctly distinguishing these relationships is essential for identifying modifiable targets for intervention and avoiding spurious causal inferences.
@@ 3.2 Data generating process
Functional principal components are calculated using the fdapace package [34]. We compare the performance of MV-FMR against U-FMR applied separately to each exposure. Performance is assessed using the Integrated Squared Error (ISE), which quantifies the average squared deviation between estimated and true effect functions, and the pointwise 95% coverage rate, defined as the proportion of time points where the 95% confidence intervals contain the true effect. Each simulation scenario is replicated $R=500$ times with a sample size $N=5000$. The number of instruments is set to $P_{1}=P_{2}=100-P12$, where $P_{12}=p_{12}\cdot 100$ with $p_{12}=0.15$ in the baseline simulations. Additional simulations vary the degree of instrument overlap across exposures, considering values of $p_{12}$ equal to 0.15, 0.3, 0.5 and 0.8, to explore how different levels of overlap affect MV-FMR performance. Instrument strength is measured using the conditional F-statistic for each principal component [22].
@@ 3.3 Simulation results
We then evaluated the robustness of MV-FMR to different functional forms of the exposure-outcome relationships for Scenarios 1 and 3. Specifically, we considered three distinct combinations: (i) both exposures with linear effects, (ii) one exposure with a linear effect and the other with a quadratic effect, and (iii) both exposures with quadratic effects. Figure 3 presents the estimated coefficient functions $\beta_{1}(t)$ and $\beta_{2}(t)$ averaged across simulations for a continuous outcome, with corresponding results for a binary outcome provided in Supplementary Figure 1. Across all time points, MV-FMR consistently achieved lower mean squared error (MSE) than the univariable approach and more accurately recovered the true functional forms. Notably, MV-FMR successfully captured nonlinear time-varying effects, whereas the univariable failed to reproduce the underlying functional relationships. Additional sensitivity analyses examined stepwise effects and decreasing lifetime effects, with MV-FMR maintaining superior performance across all specifications are shown in Supplementary Table 3. Together, these results confirm the robustness of MV-FMR across a wider range of true effect shapes.
@@
We applied the proposed MV-FMR framework to investigate the time-varying causal effects of systolic blood pressure (SBP) and body mass index (BMI) on coronary artery disease (CAD) risk. This application provides a good case study for our functional extension of multivariable MR for several reasons. First, SBP and BMI are well-established cardiovascular risk factors with substantial phenotypic correlation [8, 6], creating a scenario in which multivariable approaches offer advantages over univariable methods. Second, biological evidence indicates that BMI influences cardiovascular outcomes both directly and indirectly through blood pressure [24], establishing a mediation pathway that traditional univariable MR cannot disentangle. Indeed, cross-sectional multivariable MR studies have shown that approximately 27% of the BMI-CAD association operates through SBP [5], providing a well-characterized benchmark for validation. Third, both exposures exhibit time-varying effects over the lifecourse, with evidence suggesting that critical exposure windows for cardiovascular risk may differ between midlife and later life [7, 3]. By jointly modeling SBP and BMI trajectories while accounting for their correlation and potential mediation, MV-FMR enables estimation of the direct time-varying causal effect of each exposure on CAD risk, allowing us to recover established causal patterns and uncover critical windows of effect that cannot be identified using standard cross-sectional or univariable approaches.
@@
To address this question, we used data from UK Biobank, a large-scale prospective cohort study that recruited over 500,000 participants aged 40-69 years between 2006 and 2010 across the United Kingdom [27]. For this analysis, we restricted to participants who met the following criteria: (i) European ancestry, (ii) availability of genetic data, and (iii) at least two longitudinal measurements of both SBP and BMI between ages 50 and 70 years and before the CAD incidence. After applying these criteria, 59,907 participants were included in the final sample. SBP and BMI trajectories were reconstructed from all available measurements in primary care records and UK Biobank baseline assessment visits within the 50-70 year window. Incident CAD was defined using first-occurrence ICD-10 codes I21-I25, covering acute myocardial infarction and chronic ischemic heart disease.
@@
We investigated the effects of SBP and BMI on CAD using 274 and 199 genetic variants, respectively. Genetic instruments were obtained from large-scale genome-wide association study (GWAS) meta-analyses: for SBP were taken from Evangelou et al. [4], and for BMI from Winkler et al. [31]. Five variants were associated with both exposures, representing overlapping instruments that may introduce pleiotropy between SBP and BMI. We then applied MV-FMR to jointly estimate the time-varying causal effects of SBP and BMI on CAD risk over ages 50-70 years. Individual exposure trajectories were modeled using FPCA with PACE, with the number of components chosen via our data-driven k-fold cross-validation procedure. Two principal components were selected for both exposures. For SBP, the first two components explained 77.0% and 16.5% of the variability, respectively, while for BMI, they explained 90.5% and 8.1%. The large proportion of explained variance indicates that trajectories of both exposures were well captured by the two-component FPCA representation. Instrument strength was assessed using F-statistics for each principal component. For SBP, the first and second components yielded conditional F-statistics of 9.77 and 3.44, respectively, while for BMI, the corresponding values were 9.03 and 2.57. Although components fell below the conventional threshold for strong instruments ($F<10$), our simulation results indicate that functional MR approaches remain unbiased in such settings. To assess the benefits of joint modeling, we compared MV-FMR with U-FMR applied separately to each exposure. This comparison allowed us to evaluate whether simultaneous modeling of SBP and BMI revealed different causal effect patterns compared to the univariable analyses. Figure 4 presents the MV-FMR estimates. The analysis revealed distinct time-varying causal effects of SBP and BMI on CAD risk. Both exposures showed positive associations with CAD risk between ages 50 and 60, but their effects attenuated thereafter, becoming non-significant (95% bootstrap CI) beyond age 60. These findings suggest that the critical exposure window for SBP and BMI in relation to CAD risk is concentrated in the sixth decade of life.
@@
The comparison between MV-FMR and U-FMR provides evidence consistent with established mediation patterns. For SBP, the multivariable and univariable analyses yielded broadly consistent results, both identifying similar effect magnitudes and time windows of significance. For BMI, MV-FMR estimates showed attenuation in the point estimates compared to the univariable approach, suggesting that part of BMI’s apparent effect on CAD operates indirectly through blood pressure pathways. Inspection of the confidence intervals reveals partial overlap between MV-FMR and U-FMR estimates for BMI. Notably, this overlapping is also observed in the traditional MV-MR analysis [5], where the univariable BMI effect (OR 1.49, 95% CI 1.39 to 1.60) and the multivariable effect (OR 1.34, 95% CI 1.24 to 1.45) showed overlapping confidence intervals.